# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 5.1.2. Independent of the choice of the polarization
such that $\mathcal{E}_{W,i}/\mathcal{E}_{W,i-1}=\mathcal{Q}_{W,i}$, where $\mathcal{E}_{W,l}=\left((\mu_{\alpha})_{*}(\mathcal{E}_{\mathrm{orb},i})_{\alpha})\right)^{G_{\alpha}}$ and $\mathcal{E}_{\mathrm{orb},i}=(V_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb}}|_{V_{\mathrm{orb}},i})$. Lemma 5.1 implies that each $\mathcal{E}_{W,i}$ is locally free and so is $\mathcal{E}_{W}$. Moreover, the orbifold Higgs field $\theta_{\mathrm{orb}}$ descends first to a Higgs field on $\mathcal{E}_{W}|_{W_{\mathrm{reg}}}$, and then extends to a Higgs field $\theta_{W}$ on $\mathcal{E}_{W}$ by Lemma 2.4, which gives a locally free Higgs sheaf structure of the underlying $C^{\infty}$-bundle of $E_{\rho_{W}}$. Then $(V_{Y},\bar{\partial}_{Y},\theta_{Y})=\pi_{1}^{*}(\mathcal{E}_{W},\theta_{W})$ is a locally free Higgs sheaf structure of $(V_{Y},D_{Y})$.
From the above construction, each $V_{Y,i}$ is $\theta_{Y}$-invariant and $\bar{\partial}_{Y}$-invariant, and the induced Higgs structure of $(V_{Y},\bar{\partial}_{Y},\theta_{Y})$ on the quotient $Q_{Y,i}$ coincides with $(Q_{Y,i},\bar{\partial}_{Y,i},\theta_{Y,i})$ (because these hold on a Zariski open subset of $Y$). According to the explicit construction of the non-abelian Hodge correspondence for compact Kähler manifolds (cf. [93, Section 9] and [23]), we have $\mu_{Y}(V_{Y},\bar{\partial}_{Y},\theta_{Y})={E_{Y}}=\mu_{Y}(\mathcal{E}_{Y},\theta_{Y}),$ which implies $(\mathcal{E}_{X},\theta_{X})=\pi_{1}^{*}(\mathcal{E}_{W},\theta_{W})$.
Finally, let us verify that the reflexive sheaf $\mathcal{E}_{X}:=(g_{*}\mathcal{E}_{W})^{\vee\vee}$ is locally free, which completes the proof. $\theta_{W}$ induces a Higgs field of $\mathcal{E}_{X}$ in codimension $2$ and so induces a Higgs field $\theta_{X_{\mathrm{reg}}}$ of $\mathcal{E}_{X_{\mathrm{reg}}}$. By the above construction, $\mathcal{E}_{X}$ admits a filtration of reflexive sheaves
@@
where $\mathcal{E}_{X,i}:=(g_{*}\mathcal{E}_{W,i})^{\vee\vee}$. It is straightforward to see that each $\mathcal{E}_{X,i}$ is saturated in $\mathcal{E}_{X}$. This follows from the reflexivity of these sheaves and from the fact that $\mathcal{E}_{X,i}/\mathcal{E}_{X,i-1}=\mathcal{Q}_{X,i}$ outside the indeterminacy locus $S$ of $g^{-1}$. Again by reflexivity, we have
@@
and its orbifold Chern classes vanish. Recall that $S$ has codimension at least $3$; consequently, the orbifold Chern classes of each $\mathcal{E}_{X,i}$ vanish because $g_{\mathrm{orb}}^{[*]}\mathcal{E}_{X,i}$ coincides with $E_{\mathrm{orb},i}$ outside $g^{-1}(S)$ and all orbifold Chern classes of $E_{\mathrm{orb},i}$ vanish. Combining these statements with Proposition 2.15, we deduce that $\mathcal{E}_{X,i}/\mathcal{E}_{X,i-1}$ is locally free in codimension $3$. Applying Lemma 5.1 then shows that each $\mathcal{E}_{X,i}$ is locally free by induction on $i=1,\cdots,m$, and hence so is $\mathcal{E}_{X}$. ∎
@@ 5.3. Construction of the correspondence
Now let us complete the proof of Theorem 1.3 and Theorem 1.7. We will always assume that $X$ is a compact Kähler klt space.
@@
We begin with constructing the functor $\mu_{X}$. For any $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$, Theorem 5.5 implies $\pi^{*}(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{\widehat{X}}$ for a resolution of singularities $\pi:\widehat{X}\rightarrow X$. Combining the item (4.1.2), we could define $\mu_{X}$ as below.
@@
$\mu_{X}:\mathrm{Higgs}_{X}\rightarrow\mathrm{LSys}_{X}$ is defined by
@@
where $\pi$ is a resolution of singularities $\pi:\widehat{X}\rightarrow X$.
It can be easily seen that the diagram (1.2) commutes. It’s necessary to check that $\mu_{X}$ is independent of the choice of $\pi$. Let $\pi_{1}:X_{1}\rightarrow X$ and $\pi_{2}:X_{2}\rightarrow X$ be two resolution of singularities, we need to verify the following statement.
Lemma 5.10 (The well-definedness of $\mu_{X}$).
@@
Let $\widehat{X}$ be a desingularisation of the (unique) irreducible component of the fibre product $X_{1}\times_{X}X_{2}$ and so we have the following natural commuting diagram
@@
Note that
$g^{*}(\pi_{1}^{*})^{-1}\left(\mu_{X_{1}}(\pi_{1}^{*}(\mathcal{E}_{X},\theta_{X}))\right)$ $=f_{1}^{*}\left(\mu_{X_{1}}(\pi_{1}^{*}(\mathcal{E}_{X},\theta_{X}))\right)=\mu_{\widehat{X}}\left(f_{1}^{*}\pi_{1}^{*}(\mathcal{E}_{X},\theta_{X})\right)$
@@
$=g^{*}(\pi_{2}^{*})^{-1}\left(\mu_{X_{2}}(\pi_{2}^{*}(\mathcal{E}_{X},\theta_{X}))\right)$
where the second equality and the fourth equality follow from the functoriality of the non-abelian Hodge correspondence for compact Kähler manifolds. Because $g$ is also a resolution of singularities and so $g^{*}$ is an isomorphism by the item (4.1.2), which completes the proof. ∎
According to the above construction, it’s obvious that if $(\mathcal{E}_{X},\theta_{X})\in\mathrm{pHiggs}_{X}$ then $\mu_{X}(\mathcal{E}_{X},\theta_{X})\in\mathrm{sLSys}_{X}$. Also, we have the following functoriality building on the existence of functorial resolution of singularities (see e.g. [22, Theorem 3.10]) and the functoriality of $\mu_{X}$ for compact Kähler manifolds.
@@
To prove that $\mu_{X}$ is one-to-one, it suffices to construct its inverse map. Let $E_{X}\in\mathrm{LSys}_{X}$. Suppose that $\pi^{*}E_{X}\in\mathrm{LSys}_{\widehat{X}}$. It follows from Theorem 1.15 that there exists $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$ such that $\mu_{\widehat{X}}^{-1}(\pi^{*}E_{X})=\pi^{*}(\mathcal{E}_{X},\theta_{X})$.
@@
$\eta_{X}:\mathrm{LSys}_{X}\rightarrow\mathrm{Higgs}_{X}$ is defined by sending $E_{X}$ to $(\mathcal{E}_{X},\theta_{X})$ which is constructed as above.
It follows immediately from the above construction that $\eta_{X}\circ\mu_{X}=\operatorname{Id}$, $\mu_{X}\circ\eta_{X}=\operatorname{Id}$ and $\mu_{X}$ satisfies the properties list in Theorem 1.3.
@@
One could verify that $\mu_{X}$ is in fact functorial with respect to any morphism between compact Kähler klt spaces as in [38, page 318].
@@
Let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover of $X$ induced by some $G\subset\mathrm{Aut}(Y)$. For any $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{Higgs}_{X_{\mathrm{reg}}}$, $(\mathcal{E}_{Y_{\mathrm{reg}}},\theta_{Y_{\mathrm{reg}}})=\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is locally free and uniquely extends to a locally Higgs sheaf $(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{Higgs}_{Y}$ by Theorem 1.12. This also implies $\mathcal{E}_{X_{\mathrm{reg}}}$ is locally free because $\gamma$ is étale over $X_{\mathrm{reg}}$. Then $\gamma^{*}:\mathrm{Higgs}_{X_{\mathrm{reg}}}\rightarrow\mathrm{Higgs}_{Y}$ is defined by sending $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ to $(\mathcal{E}_{Y},\theta_{Y})$.
By the reflexivity, $(\mathcal{E}_{Y},\theta_{Y})$ is $G$-equivariant and thus $\mu_{Y}(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{LSys}_{X}$ is $G$-equivariant by Corollary 5.11. Recall $\gamma$ is finite and étale over $X_{\mathrm{reg}}$, $\mu_{Y}(\mathcal{E}_{Y},\theta_{Y})|_{\gamma^{-1}(X_{\mathrm{reg}})}$ descents to a local system $E_{X_{\mathrm{reg}}}\in\mathrm{LSys}_{X_{\mathrm{reg}}}$.
@@
$\mu_{X_{\mathrm{reg}}}:\mathrm{Higgs}_{X_{\mathrm{reg}}}\rightarrow\mathrm{LSys}_{X_{\mathrm{reg}}}$ is defined by
$(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\mapsto\left(\gamma^{\circ}_{*}(\mu_{Y}(\gamma^{*}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})))\right)^{G},$
where $\gamma:Y\rightarrow X$ is a maximally quasi-étale cover of $X$ and $\gamma^{\circ}=\gamma|_{\gamma^{-1}(X_{\mathrm{reg}})}$.
To illustrate that $\mu_{X_{\mathrm{reg}}}$ is well-defined, it’s necessary to check that
@@
$\mu_{X_{\mathrm{reg}}}$ is independent of the choice of maximally quasi-étale covers.
@@
Let $\gamma_{1}:Y_{1}\rightarrow X$ and $\gamma_{2}:Y_{2}\rightarrow X$ be two maximally quasi-étale covers of $X$. Then the fiber product $Y:=Y_{1}\times_{X}Y_{2}$ is a quasi-étale cover of $Y_{1}$. By Definition 4.2 (1), $Y$ is a étale cover of $Y_{1}$ and thus is a maximally quasi-étale cover of $X$. Then, to prove the statement, we may assume that $\gamma_{12}:Y_{1}\rightarrow Y_{2}$ is an étale cover. We adopt the associated symbols as in the aforementioned construction. By the construction. it suffices to check that $E_{Y_{1,\mathrm{reg}}}=\gamma_{12}^{*}E_{Y_{2,\mathrm{reg}}}$. Due to the reflexivity, we in fact have $(\mathcal{E}_{Y_{1}},\theta_{Y_{1}})=\gamma_{12}^{*}(\mathcal{E}_{Y_{2}},\theta_{Y_{2}})$. Applying Corollary 5.11, the proof is complete. ∎
To verify that $\mu_{X_{\mathrm{reg}}}$ is one-to-one, it suffices to construct its inverse map. Let $E_{X_{\mathrm{reg}}}\in\mathrm{LSys}_{X_{\mathrm{reg}}}$. The item (4.6.4) implies that $(\gamma|_{\gamma^{-1}(X_{\mathrm{reg}})})^{*}E_{X_{\mathrm{reg}}}$ uniquely extends to some $G$-equivariant local system $E_{Y}\in\mathrm{LSys}_{Y}$. Then $\gamma^{*}:\mathrm{LSys}_{X_{\mathrm{reg}}}\rightarrow\mathrm{LSys}_{Y}$ is defined by sending $E_{X_{\mathrm{reg}}}$ to $E_{Y}$. Building on Corollary 5.11 and the fact that $\mu_{Y}$ is one-to-one, $\mu_{Y}^{-1}(E_{Y})\in\mathrm{Higgs}_{Y}$ is $G$-equivariant and thus $\mu_{Y}^{-1}(E_{Y})|_{\gamma^{-1}(X_{\mathrm{reg}})}$ descend to some $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{Higgs}_{X_{\mathrm{reg}}}$ (again by the item ((4.6.3))).
@@
$E_{X_{\mathrm{reg}}}\mapsto\left(\gamma^{\circ}_{*}(\mu_{Y}^{-1}(\gamma^{*}E_{X_{\mathrm{reg}}}))\right)^{G}$
It follows immediately from the construction that $\mu_{X_{\mathrm{reg}}}\circ\eta_{X_{\mathrm{reg}}}=\operatorname{Id}$, $\eta_{X_{\mathrm{reg}}}\circ\mu_{X_{\mathrm{reg}}}=\operatorname{Id}$ and $\mu_{X_{\mathrm{reg}}}$ satisfies the remaining part of Theorem 1.7. The proof is complete.

### v2

@@ 5.1.2. Independence of the choice of Kähler polarization
such that $\mathcal{E}_{W,i}/\mathcal{E}_{W,i-1}=\mathcal{Q}_{W,i}$, where $\mathcal{E}_{W,i}=\left((\mu_{\alpha})_{*}(\mathcal{E}_{\mathrm{orb},i})_{\alpha})\right)^{G_{\alpha}}$ and $\mathcal{E}_{\mathrm{orb},i}=(V_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb}}|_{V_{\mathrm{orb},i}})$. Lemma 5.1 implies that each $\mathcal{E}_{W,i}$ is locally free, and thus so is $\mathcal{E}_{W}$. Moreover, the orbifold Higgs field $\theta_{\mathrm{orb}}$ first descends to a Higgs field on $\mathcal{E}_{W}|_{W_{\mathrm{reg}}}$, and then extends to a Higgs field $\theta_{W}$ on $\mathcal{E}_{W}$ by Lemma 2.4. This endows the underlying $C^{\infty}$-bundle of $E_{\rho_{W}}$ with a locally free Higgs sheaf structure. Consequently, $(V_{Y},\bar{\partial}_{Y},\theta_{Y})=pr_{1}^{*}(\mathcal{E}_{W},\theta_{W})$ is a locally free Higgs sheaf on $(V_{Y},D_{Y})$.
From the above construction, each $V_{Y,i}$ is $\theta_{Y}$-invariant and $\bar{\partial}_{Y}$-invariant, and the Higgs structure induced on the quotient $Q_{Y,i}$ coincides with $(Q_{Y,i},\bar{\partial}_{Y,i},\theta_{Y,i})$ (because the two agree on a Zariski open subset of $Y$). According to the explicit construction of the non-abelian Hodge correspondence for compact Kähler manifolds (cf. [95, Section 9] and [23]), we have $\mu_{Y}(V_{Y},\bar{\partial}_{Y},\theta_{Y})={E_{Y}}=\mu_{Y}(\mathcal{E}_{Y},\theta_{Y}),$ which implies $(\mathcal{E}_{Y},\theta_{Y})=pr_{1}^{*}(\mathcal{E}_{W},\theta_{W})$.
Finally, we verify that the reflexive sheaf $\mathcal{E}_{X}:=(g_{*}\mathcal{E}_{W})^{\vee\vee}$ is locally free, which will complete the proof. The Higgs field $\theta_{W}$ induces a Higgs field on $\mathcal{E}_{X}$ in codimension $2$, and hence induces a Higgs field $\theta_{X_{\mathrm{reg}}}$ on $\mathcal{E}_{X_{\mathrm{reg}}}$. From the above construction, $\mathcal{E}_{X}$ admits a filtration by reflexive subsheaves
@@
where $\mathcal{E}_{X,i}:=(g_{*}\mathcal{E}_{W,i})^{\vee\vee}$. It is straightforward to see that each $\mathcal{E}_{X,i}$ is saturated in $\mathcal{E}_{X}$. This follows from the reflexivity of these sheaves together with the fact that $\mathcal{E}_{X,i}/\mathcal{E}_{X,i-1}=\mathcal{Q}_{X,i}$ outside the indeterminacy locus $S$ of $g^{-1}$. Again by reflexivity, we have
@@
and its orbifold Chern classes vanish. Recall that $S$ has codimension at least $3$; consequently, the orbifold Chern classes of each $\mathcal{E}_{X,i}$ vanish, because $g_{\mathrm{orb}}^{[*]}\mathcal{E}_{X,i}$ coincides with $E_{\mathrm{orb},i}$ outside $g^{-1}(S)$ and all orbifold Chern classes of $E_{\mathrm{orb},i}$ vanish. Combining these facts with Proposition 2.15, we deduce that $\mathcal{E}_{X,i}/\mathcal{E}_{X,i-1}$ is locally free in codimension $3$. Applying Lemma 5.1 and proceeding by induction on $i=1,\cdots,m$ shows that each $\mathcal{E}_{X,i}$ is locally free, and hence so is $\mathcal{E}_{X}$. ∎
@@ 5.3. Construction of the correspondence
This subsection completes the proofs of Theorem 1.3 and Theorem 1.7. Assuming that $X$ is a compact Kähler klt space, we outline the construction of $\mu_{X}$ and $\mu_{X_{\mathrm{reg}}}$.
@@
We first construct the functor $\mu_{X}$. For any $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$, Theorem 5.5 implies $\pi^{*}(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{\widehat{X}}$ for a resolution of singularities $\pi:\widehat{X}\rightarrow X$. Together with item (4.1.b), this allows us to define $\mu_{X}$ as follows.
@@
The functor $\mu_{X}:\mathrm{Higgs}_{X}\rightarrow\mathrm{LSys}_{X}$ is defined by
@@
where $\pi:\widehat{X}\rightarrow X$ is a resolution of singularities.
It is clear that the diagram (1.2) commutes. It is necessary to check that $\mu_{X}$ is independent of the choice of $\pi$. Let $\pi_{1}:X_{1}\rightarrow X$ and $\pi_{2}:X_{2}\rightarrow X$ be two resolutions of singularities. We need to verify the following lemma.
Lemma 5.10 (Well-definedness of $\mu_{X}$).
@@
Let $\widehat{X}$ be a desingularization of the (unique) irreducible component of the fibre product $X_{1}\times_{X}X_{2}$. We then have a commutative diagram
@@
where $f_{1}$ and $f_{2}$ are the morphisms induced by the natural projections, and $g:=\pi_{1}\circ f_{1}=\pi_{2}\circ f_{2}$. Note that
$g^{*}\left((\pi_{1}^{*})^{-1}\left(\mu_{X_{1}}(\pi_{1}^{*}(\mathcal{E}_{X},\theta_{X}))\right)\right)$ $=f_{1}^{*}\left(\mu_{X_{1}}(\pi_{1}^{*}(\mathcal{E}_{X},\theta_{X}))\right)=\mu_{\widehat{X}}\left(f_{1}^{*}\pi_{1}^{*}(\mathcal{E}_{X},\theta_{X})\right)$
@@
$=g^{*}\left((\pi_{2}^{*})^{-1}\left(\mu_{X_{2}}(\pi_{2}^{*}(\mathcal{E}_{X},\theta_{X}))\right)\right),$
where the second and the fourth equalities follow from the functoriality of the non-abelian Hodge correspondence for compact Kähler manifolds. Because $g$ is also a resolution of singularities, the induced morphism $g^{*}:\mathrm{LSys}_{X}\rightarrow\mathrm{LSys}_{\widehat{X}}$ is one-to-one by item (4.1.b). This implies (5.14) and completes the proof. ∎
According to the above construction, it is obvious that if $(\mathcal{E}_{X},\theta_{X})\in\mathrm{pHiggs}_{X}$ then $\mu_{X}(\mathcal{E}_{X},\theta_{X})\in\mathrm{sLSys}_{X}$. Moreover, using the existence of functorial resolutions of singularities (see e.g. [22, Theorem 3.10]) and the functoriality of $\mu_{X}$ for compact Kähler manifolds, we obtain the following corollary, which will be used to construct $\mu_{X_{\mathrm{reg}}}$.
@@
To prove that $\mu_{X}$ is one-to-one, it suffices to construct its inverse. Let $E_{X}\in\mathrm{LSys}_{X}$. Suppose that $\pi:\widehat{X}\rightarrow X$ is a resolution of singularities and then $\pi^{*}E_{X}\in\mathrm{LSys}_{\widehat{X}}$. It follows from Theorem 1.15 that there exists $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$ such that $\mu_{\widehat{X}}^{-1}(\pi^{*}E_{X})=\pi^{*}(\mathcal{E}_{X},\theta_{X})$.
@@
The functor $\eta_{X}:\mathrm{LSys}_{X}\rightarrow\mathrm{Higgs}_{X}$ is defined by sending $E_{X}$ to $(\mathcal{E}_{X},\theta_{X})$ constructed above.
It follows immediately from the above construction that $\eta_{X}\circ\mu_{X}=\operatorname{Id}$, $\mu_{X}\circ\eta_{X}=\operatorname{Id}$ and $\mu_{X}$ satisfies the properties listed in Theorem 1.3.
@@
One might naturally ask whether adapting the argument in [38, Section 8.2] would allow us to show that $\eta_{X}$ (respectively $\mu_{\alpha}$) is in fact functorial with respect to arbitrary morphisms between compact Kähler klt spaces. However, a key ingredient used in that proof—the analytic analogue of [47, Corollary 1.7]—remains unproven at present, as the proof of the latter relies on the choice of a very ample divisor, a step that is not available in the analytic setting.
@@
We now construct the functor $\mu_{X_{\mathrm{reg}}}$. Let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover of $X$ induced by some subgroup $G\subset\mathrm{Aut}(Y)$. For any $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{Higgs}_{X_{\mathrm{reg}}}$, the reflexive pullback $(\mathcal{E}_{Y_{\mathrm{reg}}},\theta_{Y_{\mathrm{reg}}})=\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is locally free and uniquely extends to a locally free Higgs sheaf $(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{Higgs}_{Y}$ by Theorem 1.12. This also implies $\mathcal{E}_{X_{\mathrm{reg}}}$ itself is locally free because $\gamma$ is étale over $X_{\mathrm{reg}}$.
@@
$\gamma^{*}:\mathrm{Higgs}_{X_{\mathrm{reg}}}\rightarrow\mathrm{Higgs}_{Y}$ is defined by sending $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ to $(\mathcal{E}_{Y},\theta_{Y})$ constructed above.
By reflexivity, $(\mathcal{E}_{Y},\theta_{Y})$ is $G$-equivariant and thus the local system $\mu_{Y}(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{LSys}_{Y}$ is also $G$-equivariant by Corollary 5.11. Since $\gamma$ is finite and étale over $X_{\mathrm{reg}}$, the local system $\mu_{Y}(\mathcal{E}_{Y},\theta_{Y})|_{\gamma^{-1}(X_{\mathrm{reg}})}$ descends to a local system $E_{X_{\mathrm{reg}}}\in\mathrm{LSys}_{X_{\mathrm{reg}}}$.
Definition 5.15.
Set $\gamma^{\circ}:=\gamma|_{\gamma^{-1}(X_{\mathrm{reg}})}$. $\mu_{X_{\mathrm{reg}}}:\mathrm{Higgs}_{X_{\mathrm{reg}}}\rightarrow\mathrm{LSys}_{X_{\mathrm{reg}}}$ is defined by
$(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\mapsto\left(\gamma^{\circ}_{*}(\mu_{Y}(\gamma^{*}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})))\right)^{G}.$
To illustrate that $\mu_{X_{\mathrm{reg}}}$ is well-defined, we must verify
@@
The functor $\mu_{X_{\mathrm{reg}}}$ is independent of the choice of the maximally quasi-étale cover.
@@
Let $\gamma_{1}:Y_{1}\rightarrow X$ and $\gamma_{2}:Y_{2}\rightarrow X$ be two maximally quasi-étale covers of $X$ induced by some subgroups $G_{1}\subset\mathrm{Aut}(Y_{1})$ and $G_{2}\subset\mathrm{Aut}(Y_{2})$, respectively. Consider the normalization $Y$ of the fibre product $Y_{1}\times_{X}Y_{2}$. The two natural projections $Y\rightarrow Y_{1}$ and $Y\rightarrow Y_{2}$ are quasi-étale. According to Definition 4.2 (1), $Y$ is an étale cover of $Y_{1}$ and hence also a maximally quasi-étale cover of $X$.
Therefore, to prove the lemma, we may assume that there exists an étale morphism $\gamma_{12}:Y_{1}\rightarrow Y_{2}$. Let $\pi_{2}:\widehat{Y}_{2}\rightarrow Y_{2}$ be a resolution of singularities and $\widehat{Y}_{1}$ be a desingularization of the (unique) irreducible component of the fibre product $\widehat{Y}_{2}\times_{X}Y_{1}$. Denote by $\widehat{\gamma}_{12}:\widehat{Y}_{1}\rightarrow\widehat{Y}_{2}$ and $\pi_{1}:\widehat{Y}_{1}\rightarrow Y_{1}$ the morphisms induced by the natural projections from the fibre product. We adopt the notation from the construction above. One can see that $(\mathcal{E}_{\widehat{Y}_{1}},\theta_{\widehat{Y}_{1}})=\widehat{\gamma}^{*}_{12}(\mathcal{E}_{Y_{1}},\theta_{Y_{2}})$ by the following commutative diagram
$\widehat{Y}_{1}$$Y_{1}$$\widehat{Y}_{2}$$Y_{2}.$$\scriptstyle{\pi_{1}}$$\scriptstyle{\widehat{\gamma}_{12}}$$\scriptstyle{\gamma_{12}}$$\scriptstyle{\pi_{2}}$
Applying the functoriality of the non-abelian Hodge correspondence for compact Kähler manifolds yields $E_{\widehat{Y}_{1}}=\widehat{\gamma}_{12}^{*}E_{{\widehat{Y}}_{2}}$. This completes the proof. ∎
To show that $\mu_{X_{\mathrm{reg}}}$ is one-to-one, it is enough to construct its inverse. Let $E_{X_{\mathrm{reg}}}\in\mathrm{LSys}_{X_{\mathrm{reg}}}$. Item (4.6.d) implies that $(\gamma|_{\gamma^{-1}(X_{\mathrm{reg}})})^{*}E_{X_{\mathrm{reg}}}$ uniquely extends to a $G$-equivariant local system $E_{Y}\in\mathrm{LSys}_{Y}$.
Definition 5.17.
We define $\gamma^{*}:\mathrm{LSys}_{X_{\mathrm{reg}}}\rightarrow\mathrm{LSys}_{Y}$ by sending $E_{X_{\mathrm{reg}}}$ to $E_{Y}$.
Building on Corollary 5.11 and the fact that $\mu_{Y}$ is one-to-one, the Higgs sheaf $\mu_{Y}^{-1}(E_{Y})\in\mathrm{Higgs}_{Y}$ is $G$-equivariant and thus $\mu_{Y}^{-1}(E_{Y})|_{\gamma^{-1}(X_{\mathrm{reg}})}$ descends to some $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{Higgs}_{X_{\mathrm{reg}}}$ (again by item (4.6.c)).
@@
$E_{X_{\mathrm{reg}}}\mapsto\left(\gamma^{\circ}_{*}(\mu_{Y}^{-1}(\gamma^{*}E_{X_{\mathrm{reg}}}))\right)^{G}.$
It follows immediately from the construction that $\mu_{X_{\mathrm{reg}}}\circ\eta_{X_{\mathrm{reg}}}=\operatorname{Id}$, $\eta_{X_{\mathrm{reg}}}\circ\mu_{X_{\mathrm{reg}}}=\operatorname{Id}$ and $\mu_{X_{\mathrm{reg}}}$ satisfies the remaining properties stated in Theorem 1.7. The proof is complete.